40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13789     Entry time: Wed Apr 25 19:09:37 2018     In reply to: 13786     Reply to this: 13791
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: New look EX Fiber coupling 

Summary:

I implemented most of the things outlined in my previous elog. Implementing the a la mode solution after including all lenses, I managed to achieve >90% mode-matching into the fiber. Power monitor PD has not been re-installed yet, neither has the bracket I removed. The polarization monitoring setup on the PSL table has now been hooked up to the EX fiber, let's see how it does overnight. All quoted power measurements in this elog were made with the Ophir power meter (filter off).

Details:  

Attachment #1 shows the implemented MM solution. I did not include the PBS substrate in the calculation, maybe that will help a little.

Attachment #2 shows the new layout. The beam is a little low on the PBS and HWP - I will swap these out to mounts with slightly lower height, that should improve the situation a little. There is no evidence of clipping, and the beam clears all edges by at least 3 beam diameters.

Attachments #3 and #4 show the EX fiber before and after cleaning respectively. Seems like the cleaning was successful.

Attachment #5 shows the beam incident on the coupler with on an IR card. This beam only goes through a QWP, lens, BS and 45 degree steering mirror, so I'm not sure what's responsible for the large halo around the main beam. There is significant power in the halo too - I measured 25mW right before the coupler, but if I use an iris to try and cut off the halo, the power is measured to be ~19mW.

Alignment Procedure:

  • Connect spare fiber such that I can monitor coupled power (minus fiber losses and joint loss) at EX table.
  • Use Fluke fault analyzer to align input and collimator modes coarsely.
  • Monitored coupled power continuously using Fiber Power Meter (although MM calculations were made with Ophir, this was more convenient for "Live" viewing).
  • Tweaked one available steering mirror and K6XS axes to maximize coupled power. 
  • Tweaked lens positions slightly to see if significant improvement could be made.
  • After optimizing, I measured 17.1mW coming out of the EX fiber at the PSL table. As mentioned earlier, the input power is tricky to measure given the large amount of junk light around the main mode. But I measured 18.6 mW after the iris. So this is ~95%. In any case, safe to say that we are waaaay better than the previous situation of 380uW out of 1.9mW. 
  • Added PBS and HWP to cut the incident power to 1.6mW. I measured 1.2mW on the PSL table. Probably adding the PBS screwed up the MM a bit, to be tweaked tomorrow. 
  • I had moved the Green shutter a bit during this work - as a result, the Green REFL was not making it back to the REFL PD. I remedied this, and EX Green TEM00 mode was locked to the arm. GTRX of ~0.4 was recovered, which is around the number I'm used to seeing.
Attachment 1: EX_fiber_MM.pdf  71 kB  Uploaded Wed Apr 25 21:18:14 2018  | Hide | Hide all
EX_fiber_MM.pdf
Attachment 2: IMG_6977.JPG  2.249 MB  Uploaded Wed Apr 25 21:19:58 2018  | Hide | Hide all
IMG_6977.JPG
Attachment 3: IMG_6972.JPG  1.174 MB  Uploaded Wed Apr 25 21:20:13 2018  | Hide | Hide all
IMG_6972.JPG
Attachment 4: IMG_6974.JPG  1.034 MB  Uploaded Wed Apr 25 21:20:22 2018  | Hide | Hide all
IMG_6974.JPG
Attachment 5: IMG_6976.JPG  1.197 MB  Uploaded Wed Apr 25 21:20:40 2018  | Hide | Hide all
IMG_6976.JPG
ELOG V3.1.3-